# Pulse-induced memory-like effect in cyclotron motion?

### v1

@@ Abstract
We study how a charged particle moving in a uniform magnetic field along its standard circular path (cyclotron motion) reacts to a short-duration, homogeneous, uniform electric field pulse injected in the plane perpendicular to the magnetic field. A ‘permanent’ change in the radius of the initial circle and a shift of its centre is noted at later times, after the pulse is switched off. The magnitude of the velocity undergoes a change too, akin to a ‘velocity kick’. In summary, our results suggest a pulse-induced ‘electromagnetic memory-like effect’, which is not quite a ‘wave memory’, but, nevertheless, has similar features within a simple, non-relativistic context.
@@ I Introduction
where ${\bf E}$ is the electric field, usually taken as a radiation field due to a source far away. In the above formula, if one considers an electric field which is non-zero and constant only over a small interval (0,T) in time, then the R. H. S. simply becomes ${q\bf E_{0}}T$ and we have the velocity at negative infinity (past) having a different value compared to its value at positive infinity (future). This also implies a jump or a kick in the velocity, imparted to the charged test particle by the electric pulse. Conventionally though, one considers the electric field due to radiation–say, electric dipole radiation (far away from the source) – and its effect on a test charge. In the non-relativistic case, one can evaluate the time rate of change of the dipole moment of the source in the past and future and relate it to the charges and velocities of the source. This would give an expression for the velocity kick on the test charge in terms of the behaviour of the source in the past and future biegarf2 . The analysis is somewhat more intricate when one moves away from the slow motion approximation for the source entities. It is available in biegarf2 . The essence however is to see if the velocity of the test charge changes on account of the field or the pulse (i.e. we find a nonzero value of the L. H. S. in Eqn. (1)), and whether there is a ‘permanent change’ which appears after the pulse has departed, and persists later too. It is this viewpoint, which does not necessarily attribute memory to a passing wave field, is what we focus on in this article. Diverse aspects of electromagnetic memory as well as ‘color’ (Yang-Mills) memory have been discussed in numerous papers in the recent past winicour ; pasterski ; pate ; satish ; andrea ; garfinkle ; biegarf2 ; seraj .
In the example which we discuss here, we will consider a particle of charge $q$ and mass $m$ placed in a uniform magnetic field and encountering a spatially homogeneous electric pulse. Prior to the arrival of the pulse, the particle travels in a circular path. After the pulse has departed, what happens to this trajectory? How is it influenced by the pulse? Is that influence a permanent change? These questions when answered (as we do below) will shed light on an effect quite similar to ‘electromagnetic memory’ , albeit in a rather simple, purely non-relativistic context. We emphasize here, that we are not considering a ‘wave’ memory–rather we are defining ‘memory’ arising out of a short duration electric pulse in a background uniform magnetic field. From our perspective, the memory aspect is largely encoded in a ‘persistent and permanent change’ caused by the pulse and we are interested in quantifying this change in our work.
@@ II Charged particle in a uniform magnetic field and an electric pulse
In the standard cyclotron, the electric field is chosen to be an alternating harmonic field applied between the dees and oscillating at a frequency in the radio-frequency range. This leads to a spiral path and the charged particle gets accelerated in the process, thereby achieving the aim of an accelerator chautard . In contrast here we apply a short duration electric field with the aim of injecting a visible difference between trajectories before and after the duration of the pulse.
@@
For the problem at hand, we will need to write down solutions for $x(t)$, $y(t)$, $v_{x}(t)$, $v_{y}(t)$ in each of the three regions: I ($t\leq 0$), II($0\leq t\leq T$ and III($t\geq T$). The solutions for the coordinates and velocities have to be matched at the two boundaries at $t=0$ and $t=T$. This problem appears to have been discussed briefly in a very old paper gould , though the analysis there addresses a different motivation and purpose. We also know that the coupled equations can be rewritten using a complex velocity ${\bf\tilde{w}}=v_{x}+iv_{y}$ and solved directly using ${\bf\tilde{w}}={\bf\tilde{p}}\,e^{-i\omega_{B}t}$. The boundary conditions mentioned earlier will have to be imposed after writing down the solutions in each region.
@@
In Region I, one can clearly see that we have a circular trajectory given by the equation:
@@
The effect of the pulse is manifest in the nature of the trajectories in regions II and III. In III, we find that the trajectory is still a circle but with a shifted centre and a different radius. The equation of this new circle at times $t\geq T$ is given as:
@@
Therefore, the centre of the circle now appears to be at $\left(C_{0}+\frac{E}{B}\left(\sin\,\alpha\right)T,C^{\prime}_{0}-\frac{E}{B}\left(\cos\,\alpha\right)T\right)$. Its radius is now changed (increased/decreased) to a a value equal to the square root of the expression given in the R. H. S. of Eqn. (20). It is easy to check and worth noting that there is no choice of values of any of the constants for which both the location of the centre and the radius remain unchanged.
@@
An observer watching the charge from a time much before $t=0$ will find it moving along a circular path of a certain radius. As one approaches the time $t=0$, the charge jumps to the trajectory (not closed) as specified in Eqns. (11) and (12). When $t=T$ is reached, the trajectory joins up with a circle with a different centre and a different radius. Subsequently, the charge is found to move on this new circular path forever. Figure 4 shows the evolution – the initial circle in red, the curved path in-between in blue and the final circular path in green. The shift of the centre of the circle and the change in the radius are once again manifest here, clearly. One may extend the time to values before $t=-7$ units and beyond $t=15$ units to any extent – in such a case we will just observe winding around the red and green circles along which the charge moves in the remote past and distant future, respectively.
Let us now note a few features and consequences of the main results stated above.
$\bullet$ With properly tuned initial conditions one may end up with $x,y$ as constant. This will happen when the values of $A_{0}$ and $D_{0}$ are chosen to be numerically given by:
@@
$\bullet$ For special values of $\omega_{B}T$ we have the following scenarios. When $\omega_{B}T=2n\pi$ ($n=0,1,2...$), the radius of the circle does not change, though its centre shifts. If, $\omega_{B}T=(2n+1)\pi$ ($n=0,1,2,...$), the radius changes according to the chosen value of $\alpha$. For $\omega_{B}T=-2\alpha$, the radius always decreases.
$\bullet$ The change in the velocity components can be found by simply evaluating $\Delta v_{\alpha}=v_{\alpha}(t_{f})-v_{\alpha}(t_{i})$, where $t_{i}<0$ and $t_{f}>T$ and $\alpha\equiv x,y$. For no pair of values of $t_{i}$ and $t_{f}$ is the difference $\Delta v_{\alpha}=0$. This ensures the presence of a memory-like effect.
$\bullet$ One can also evaluate $v^{2}$ at any time $t_{i}<0$ and $t_{f}>0$. The difference $v_{f}^{2}-v_{i}^{2}$ turns out to be
$\bullet$ Let us now consider a pair of identical charges starting out with different initial conditions. We make the following choices for the constants:
@@
From the expressions above it is clear that $\Delta R_{i}\neq\Delta R_{f}$ and hence $\Delta v_{i}\neq\Delta v_{f}$. In addition the centre also shifts. This is shown in Figure 6. It also demonstrates the ‘permanent change’ character of the memory effect.
$\bullet$ The velocity field before and after the pulse have ${\bf\nabla\cdot v=0}$ and ×${\bf\nabla\times v\neq 0}$. Hence, one would say that the gradient of the velocity field has zero expansion and shear and nonzero rotation both before and after the electric pulse acts. The rotation is just proportional to $\omega_{B}$. One cannot therefore derive any link to a memory effect from the kinematics of a family of trajectories skrsadg . In other words there is no ${\cal B}$ memory olough , ic1 to be seen here.
$\bullet$ It may be possible to see the above-analysed memory effect in a cyclotron. Recall that one uses a radio-frequency alternating electric field (equal to $\omega_{B}$ and spread over a small region along the diameter of the dees) between the dees of a cyclotron, in order to drive the electron along a spiral path in the dees. In this process, the electron is also accelerated, which is indeed the purpose in a cyclotron. One will have to modify the set-up a bit so that one has a homogeneous, short-duration electric field apart from the magnets which give the uniform magnetic field. It may then be possible to see the change in the circular trajectory, caused by the pulse – i.e. see the change in radius, the shift of the centre and the change in the velocity. The practical implementation of this idea in a possible experiment will surely require more detailed thought which is left for the interested experimentalist to worry about.
$\bullet$ In order to get a better picture, let us now find some numbers. We assume $m=3.3\times{10}^{-27}$ kg (deuteron) , $q=1.6\times{10}^{-19}$ C and $B=1.5$ Tesla. $\alpha$ is chosen to be equal to $2\pi$. These values imply $\omega_{B}=72.7\times{10}^{6}$ rad/s which is in the radio frequency range (around 12 MHz). We also assume $T=10$ nanoseconds and $E=50$ MegaVolts/m. The constants are chosen as $A_{0}=0.30$ m, $D_{0}=0.40$ m, $C_{0}=0.60$ m and $C^{\prime}_{0}=-0.10$ m. The chosen values yield the following:
@@
$\bullet$ It is possible to write down the current ${\bf J}$ which produces the electric pulse using the Ampere-Maxwell law. The electric pulse can be written as
@@
$\bullet$ One may write down a formula analogous to Eqn. (1) for the system discussed here. Recall the complex velocity ${\bf\tilde{w}}=v_{x}+iv_{y}={\bf\tilde{p}}\,e^{-i\omega_{B}t}$. Solving the differential equation for ${\bf\tilde{w}}$, we can formally write
[…]
One can find the ‘velocity kick’ from this expression and it generalises Eqn. (1) when both electric and magnetic fields are present. In principle, it is possible to replace the electric field in terms of the current density vector (integral over it) and write down the ‘velocity kick’ solely in terms of the source properties, as has been the usual practice in memory calculations.
III Conclusions

### v2

@@ Abstract
We study how a charged particle moving in a uniform magnetic field along its standard circular path (cyclotron motion) reacts to a short-duration, homogeneous, uniform electric field pulse injected in the plane perpendicular to the magnetic field. A ‘permanent’ change in the radius of the initial circle and a shift of its centre is noted at later times, after the pulse is switched off. The velocity vector (components and magnitude) undergoes a change too, akin to a ‘velocity kick’. The cause behind this permanent change appears to be linked to the difference between the vector potentials before and after the duration of the electric pulse. Further, we show how such vector potentials in the past and future are related through a gauge transformation. In summary, our results suggest a pulse-induced ‘electromagnetic memory-like effect’, which is not quite a ‘wave memory’, but, nevertheless, has similar features within a simple, non-relativistic context.
@@ I Introduction
where ${\bf E}$ is the electric field, usually taken as a radiation field due to a source far away. In the above formula, if one considers an electric field which is non-zero and constant only over a small interval (0,T) in time, then the R. H. S. simply becomes ${q\bf E_{0}}T$ and we have the velocity at negative infinity (past) having a different value compared to its value at positive infinity (future). This also implies a jump or a kick in the velocity, imparted to the charged test particle by the electric pulse. Conventionally though, one considers the electric field due to radiation–say, electric dipole radiation (far away from the source) – and its effect on a test charge. In the non-relativistic case, one can evaluate the time rate of change of the dipole moment of the source in the past and future and relate it to the charges and velocities of the source. This would give an expression for the velocity kick on the test charge in terms of the behaviour of the source in the past and future biegarf1 . The analysis is somewhat more intricate when one moves away from the slow motion approximation for the source entities. In such a fully relativistic scenario one uses the outgoing null coordinate $u$ instead of time $t$ and evaluations are done at future (past) null infinity ($r\rightarrow\infty$) using $u$. The discussion is available in biegarf1 and also nicely summarised briefly in jokela . The essence however is to see if the velocity of the test charge changes on account of the field or the pulse (i.e. we find a nonzero value of the L. H. S. in Eqn. (1)), and whether there is a ‘permanent change’ which appears after the pulse has departed, and persists later too. It is this viewpoint, which does not necessarily attribute memory to a passing wave field at null infinity, is what we focus on in this article. In a sense, our approach to ‘memory’ as an observable phenomenon is somewhat similar to the idea of ‘persistent observables’ extensively pursued in eanna1 ; eanna2 ; eanna3 . Diverse aspects of electromagnetic memory as well as ‘color’ (Yang-Mills) memory have been discussed in numerous papers in the recent past winicour ; susskind ; pasterski ; pate ; satish ; andrea ; garfinkle ; biegarf2 ; jokela ; seraj .
In the example which we discuss here, we will consider a particle of charge $q$ and mass $m$ placed in a uniform magnetic field and encountering a spatially homogeneous electric pulse. Prior to the arrival of the pulse, the particle travels in a circular path. After the pulse has departed, what happens to this trajectory? How is it influenced by the pulse? Is that influence a permanent change? These questions when answered (as we do below) will shed light on an effect quite similar to ‘electromagnetic memory’ , albeit in a rather simple, purely non-relativistic context. We emphasize here, that we are not considering a ‘wave’ memory–rather we are defining ‘memory’ as a remembrance arising out of a short duration electric pulse in a background uniform magnetic field. From our perspective, the memory aspect is largely encoded in a ‘persistent and permanent change’ caused by the pulse and we are interested in illustrating and quantifying this change through our work.
@@ II Charged particle in a uniform magnetic field and an electric pulse
In the standard cyclotron, the electric field is chosen to be an alternating harmonic field applied between the dees and oscillating at a frequency in the radio-frequency range. This leads to a spiral path and the charged particle gets accelerated in the process, thereby achieving the aim of an accelerator chautard . In contrast, here we apply a short duration electric field with the aim of injecting a visible difference between trajectories before and after the duration of the pulse.
@@
For the problem at hand, we will need to write down solutions for $x(t)$, $y(t)$, $v_{x}(t)$, $v_{y}(t)$ in each of the three regions: I ($t\leq 0$), II($0\leq t\leq T$ and III($t\geq T$). The solutions for the coordinates and velocities have to be matched at the two boundaries at $t=0$ and $t=T$. This problem appears to have been discussed briefly in a very old paper gould , though the analysis there addresses a different motivation and purpose.
@@
We also know that the coupled equations for $v_{x}$, $v_{y}$ can be rewritten using a complex velocity ${\bf\tilde{w}}=v_{x}+iv_{y}$ and solved directly using ${\bf\tilde{w}}={\bf\tilde{p}}\,e^{-i\omega_{B}t}$. The differential equation for the complex ${\bf\tilde{p}}$ in Region II is,
$\frac{d{\bf\tilde{p}}}{dt}=\frac{qE}{m}e^{i(\omega_{B}t+\alpha)}$ (19)
In Regions I and III, the corresponding equations are found by just putting $E=0$.. Solving the differential equation for ${\bf\tilde{p}}$ in each region and using the boundary matching for ${\bf\tilde{p}}$ at $t=0$ and $t=T$, the difference between the complex velocities in Regions I and III can be found. This is given as,
${\bf\tilde{w}}_{(t>T)}-{\bf\tilde{w}}_{(t<0)}=-i\frac{E}{B}e^{-i\omega_{B}t}\left[e^{i(\omega_{B}T+\alpha)}-e^{i\alpha}\right]$ (20)
The velocity kick for the $x$ and $y$ components of ${\bf v}$ as noted from (9), (10), (17), (18) matches with the real and imaginary parts of (20). In a way, ((20) generalises Eqn. (1) when both electric and magnetic fields (uniform) are present. One may consider $t$ to be arbitrarily large ($t\rightarrow\infty$, distant future) or small ($t\rightarrow-\infty$, remote past). Subsequently, the $x(t)$ and $y(t)$ can also be found by integrating ${\bf w}=\frac{d{\bf z}}{dt}=\dot{x}+i\dot{y}$ and applying the boundary conditions. Later, we will obtain a different and perhaps more appropriate way of expressing the velocity kick, once we elaborate on the vector potential in the next section.
@@
Let us now turn to analysing the nature and features of the trajectories in the different regions. In Region I, one can clearly see that we have a circular trajectory given by the equation:
@@
The effect of the pulse is manifest in the modified trajectories in regions II and III. In III, we find that the trajectory is still a circle but with a shifted centre and a different radius. The equation of this new circle at times $t\geq T$ is given as:
@@
Therefore, the centre of the circle now appears to be at $\left(C_{0}+\frac{E}{B}\left(\sin\,\alpha\right)T,C^{\prime}_{0}-\frac{E}{B}\left(\cos\,\alpha\right)T\right)$. Its radius is now changed (increased/decreased) to a value equal to the square root of the expression given in the R. H. S. of Eqn. (22). It is easy to check and worth noting that there is no choice of values of any of the constants for which both the location of the centre and the radius remain unchanged.
@@
An observer watching the charge from a time much before $t=0$ will find it moving along a circular path of a certain radius. As one approaches the time $t=0$, the charge jumps to the trajectory (not closed) as specified in Eqns. (11) and (12). When $t=T$ is reached, the trajectory joins up with a circle with a different centre and a different radius. Thereafter, the charge is found to move on this new circular path forever. Figure 4 shows the evolution – the initial circle in red, the curved path in-between in blue and the final circular path in green. The shift of the centre of the circle and the change in the radius are once again manifest here, clearly. One may extend the time to values before $t=-7$ units and beyond $t=15$ units (also note, these are just values chosen for plotting the graphs) to any extent – in such a case we will just observe winding around the red and green circles along which the charge moves in the remote past and distant future, respectively.
We now list a few features and consequences of the equations and results stated above. The main purpose is to point out specific changes caused by the pulse which highlight memory-like characteristics.
Velocities: The change in the velocity components can be found by simply evaluating $\Delta v_{\alpha}=v_{\alpha}(t_{f})-v_{\alpha}(t_{i})$, where $t_{i}<0$ and $t_{f}>T$ and $\alpha\equiv x,y$. For no pair of values of $t_{i}$ and $t_{f}$ is the difference $\Delta v_{\alpha}=0$. We can stretch $t_{i}$ to negative infinity and $t_{f}$ to positive infinity, thereby reaching asymptotic past and future. The nonzero $\Delta v_{\alpha}$ ensures the presence of a memory-like effect. We will understand this aspect in more detail in the next section.
One can also evaluate $v^{2}$ at any time $t_{i}<0$ and $t_{f}>0$. The difference $v_{f}^{2}-v_{i}^{2}$ turns out to be
The velocity fields before and after the pulse have ${\bf\nabla\cdot v=0}$ and ×${\bf\nabla\times v\neq 0}$. Hence, one would say that the gradient of the velocity field has zero expansion and shear but a nonzero rotation both before and after the electric pulse acts. The rotation is just proportional to $\omega_{B}$. One cannot therefore derive any link to a memory effect from the kinematics of a family of trajectories skrsadg . In other words there is no ${\cal B}$ memory olough , ic1 to be seen here.
Special cases: With properly tuned initial conditions one may end up with $x,y$ as constant. This will happen when the values of $A_{0}$ and $D_{0}$ are chosen to be numerically given by:
@@
In addition, for certain chosen values of $\omega_{B}T$ we have the following scenarios. When $\omega_{B}T=2n\pi$ ($n=0,1,2...$), the radius of the circle does not change, though its centre shifts. If, $\omega_{B}T=(2n+1)\pi$ ($n=0,1,2,...$), the radius changes according to the chosen value of $\alpha$. For $\omega_{B}T=-2\alpha$, the radius always decreases as long as $0<\alpha<\pi$.
Relative separation: What happens when we consider a pair of identical charges starting out with different initial conditions? To analyse the trajectories, we make the following choices for the constants:
@@
From the expressions above it is clear that $\Delta R_{i}\neq\Delta R_{f}$ and hence $\Delta v_{i}\neq\Delta v_{f}$. Further, the centre also shifts. This is shown in Figure 6. which also demonstrates a ‘permanent change’ character of the memory-like effect. Note that in articles on gravitational wave memory (eg. bragri ,zhang1 , ic1 , ic4 , eanna1 , eanna2 ) one usually analyses the change in ‘relative separation’ between pairs of test particles, which is qualitatively analogous to studying geodesic deviation. In our discussion above, we have, in a similar way, also shown the presence of a net change in the relative separation.
Experimental possibility: It may be possible to see the above-analysed memory effect in a cyclotron. Recall that one uses a radio-frequency alternating electric field (equal to $\omega_{B}$ and spread over a small region along the diameter of the dees) between the dees of a cyclotron, in order to drive the electron along a spiral path in the dees. In this process, the electron is also accelerated, which is indeed the purpose in a cyclotron. One will have to modify the set-up a bit so that one has a homogeneous, short-duration electric field apart from the magnets which give the uniform magnetic field. It may then be possible to see the change in the circular trajectory, caused by the pulse – i.e. see the change in radius, the shift of the centre and the change in the velocity. The practical implementation of this idea in a possible experiment will surely require more detailed thought which is left for the interested experimentalist to worry about.
In order to get a realistic picture, let us now find some numbers. We assume $m=3.3\times{10}^{-27}$ kg (deuteron) , $q=1.6\times{10}^{-19}$ C and $B=1.5$ Tesla. $\alpha$ is chosen to be equal to $2\pi$. These values imply $\omega_{B}=72.7\times{10}^{6}$ rad/s which is in the radio frequency range (around 12 MHz). We also assume $T=10$ nanoseconds and $E=50$ MegaVolts/m. The constants are chosen as $A_{0}=0.30$ m, $D_{0}=0.40$ m, $C_{0}=0.60$ m and $C^{\prime}_{0}=-0.10$ m. The chosen values yield the following:
@@
Current source: One may write down the current ${\bf J}$ which produces the electric pulse using the Ampere-Maxwell law. The electric pulse is written as
@@
III THE ROLE OF THE VECTOR POTENTIAL
In the calculation presented above, we have seen a permanent change in the trajectory of the charged particle after the pulse is switched off. We have tried to suggest that this is indeed similar to ‘memory’ as understood in radiative contexts involving gravitational or electromagnetic waves. Here we provide a somewhat stronger argument which, hopefully, will bring our results closer to what is usually called ‘memory’.
Firstly, as stated at the outset, one must note that here we do not have the standard ‘null’ coordinate $u$ used in memory calculations, since we are in a non-relativistic regime. $u$ here is simply replaced by ordinary time $t$.
In our example in the previous sections we used an electric field pulse ${\bf E}$ and a uniform ${\bf B}$. The vector potential $\bf A({\bf r},t)$ which generates the chosen $\bf E$ and $\bf B$ may be written as:
${\bf A}=-\frac{1}{2}B\left(y{\bf\hat{i}}-x{\bf\hat{j}}\right)+\left(\cos\alpha{\bf\hat{i}}+\sin\alpha{\bf\hat{j}}\right)F(t)$ (30)
where,
$F(t)=0\hskip 28.90755ptt\leq 0$ (31)
$=-E\,t\hskip 14.45377pt0\leq t\leq T$ (32)
$=-E\,T\hskip 28.90755ptt\geq T$ (33)
The first term in (30), which is purely spatial, generates the uniform ${\bf B}=\nabla\times{\bf A}$ i.e. a constant magnetic field along the $z$ direction, present for all time. The second term in (30), which we will refer to as ${\bf A^{(t)}}$, is responsible for the square pulse electric field, as evident from the defining relation ${\bf E}=-\frac{\partial{\bf A}}{\partial t}=-\frac{d{\bf A^{(t)}}}{dt}$. It is worth noting here that the vector potential is different in the regions $t<0$ and $t>T$ though the $\bf E$, $\bf B$ fields remain the same, before and after the duration of the pulse. Figure 7 (the blue curve) shows the temporal part of the vector potential as a function of time along with the electric field pulse (yellow curve). The difference $\Delta\bf A^{(t)}$ between the temporal parts of the vector potentials before and after the duration of the pulse is a non-zero constant vector, given by,
$\Delta{\bf A^{(t)}}=-ET\left(cos\alpha\,{\bf\hat{i}}+\sin\alpha\,{\bf\hat{j}}\right)$ (34)
It is this change which manifests itself in the differences in the circular trajectories (shift of the centre, change in radius and velocity) and is responsible for a memory-like effect.
Figure 7: The electric field pulse (yellow curve) for $\alpha=0$, $E=1$ unit and $T=2$ units. The temporal part of the vector potential (blue curve) for $\alpha=0$, $E=1$ unit and $T=2$ units.
More importantly, we may relate the vector potential before and after the pulse via a gauge transformation. Since ${\bf A_{before}}$ is related to ${\bf A_{after}}$ via ${\bf A_{after}}={\bf A_{before}}+{\nabla\Lambda}$, we obtain,
$\Lambda=-ET\left(x\cos\alpha+y\sin\alpha\right)$ (35)
Hence, it should be clear that our memory-like effect has its roots in the gauge-related change in the value (asymptotic values in the past and future) of the vector potential ${\bf A}$. Even though the change is a constant shift, it affects motion. The fact that such a gauge transformation between vector potentials in the asymptotic past and future is associated with memory has been noted by others too susskind ; madwin2 .
In the special case where the velocity kick is given by the time integral of the electric field vector only (i.e. no magnetic field or scalar potential are present) one can clearly see that it is equal to the difference in the asymptotic values of the vector potential. More precisely, we have the positions and velocities in the three regions given as:
Region I ($t\leq 0$):
$x^{I}=v_{x0}t+x_{0}\hskip 14.45377pt;\hskip 14.45377ptv_{x}^{I}=v_{x0}\hskip 14.45377pt;\hskip 14.45377pty^{I}=v_{y0}t+y_{0}\hskip 14.45377pt;\hskip 14.45377ptv_{y}^{I}=v_{y0}$ (36)
Region II ($0\leq t\leq T$):
$x^{II}=x_{0}+v_{x0}t+\left(\frac{qE\cos\,\alpha}{m}\right)\frac{t^{2}}{2}\hskip 14.45377pt;\hskip 14.45377ptv_{x}^{II}=v_{x0}+\left(\frac{qE\cos\,\alpha}{m}\right)t$ (37)
$y^{II}=y_{0}+v_{y0}t+\left(\frac{qE\sin\,\alpha}{m}\right)\frac{t^{2}}{2}\hskip 14.45377pt;\hskip 14.45377ptv_{y}^{II}=v_{y0}+\left(\frac{qE\sin\,\alpha}{m}\right)t$ (38)
Region III ($t\geq T$):
$x^{III}=x_{0}+v_{x0}t+\left(\frac{qE\cos\,\alpha}{m}\right)T\,t-\left(\frac{qE\cos\,\alpha}{m}\right)\frac{T^{2}}{2}\hskip 8.67204pt;\hskip 8.67204ptv_{x}^{III}=v_{x0}+\left(\frac{qE\cos\,\alpha}{m}\right)T$ (39)
$y^{III}=y_{0}+v_{y0}t+\left(\frac{qE\sin\,\alpha}{m}\right)T\,t-\left(\frac{qE\sin\,\alpha}{m}\right)\frac{T^{2}}{2}\hskip 8.67204pt;\hskip 8.67204ptv_{y}^{III}=v_{y0}+\left(\frac{qE\sin\,\alpha}{m}\right)T$ (40)
It is easy to verify that $\Delta{\bf v}=-\frac{q}{m}\Delta{\bf A}=-\frac{q}{m}\nabla\Lambda$ ($\Delta{\bf v}={\bf v}(t>T)-{\bf v}(t<0)$, $\Delta{\bf A}={\bf A}(t>T)-{\bf A}(t<0)$), using the expressions in (34), (35). One also notices that for a pair of test charges with different conditions at $t=0$, there is no change in relative separation.
However, the kick is not directly given by just a difference of asymptotic values of ${\bf A}$, when we have an ambient magnetic field (uniform or non-uniform). Integrating Eqns. (4), (5) from large negative to large positive values of $t$, one obtains,
$\Delta{\bf v}=-\frac{q}{m}{\Delta}{\bf A^{(t)}}+\omega_{B}\Delta{\bf\tilde{r}}$ (41)
where $\Delta{\bf v}={\bf v}(t>T)-{\bf v}(t<0)$, $\Delta{\bf A^{(t)}}={\bf A^{(t)}}(t>T)-{\bf A^{(t)}}(t<0)$, ${\bf\tilde{r}}=y{\bf\hat{i}}-x{\bf\hat{j}}$ and $\Delta{\bf\tilde{r}}={\bf\tilde{r}}(t>T)-{\bf\tilde{r}}(t<0)$. Equivalently, one may use the total vector potential ${\bf A}$ instead of $\bf A^{(t)}$ in (41), though one needs to keep in mind that $\frac{d{\bf A}}{dt}=\frac{\partial{\bf A}}{\partial t}+\left({\bf v}\cdot\nabla\right){\bf A}=\frac{\partial{\bf A}}{\partial t}-\frac{B}{2}\frac{d{\bf\tilde{r}}}{dt}$ and, therefore, $\Delta{\bf A^{(t)}}=\Delta{\bf A}+\frac{B}{2}{\Delta{\bf\tilde{r}}}$. Thus, we may rewrite (41) as: $\Delta{\bf v}=-\frac{q}{m}{\Delta}{\bf A}+\frac{\omega_{B}}{2}\Delta{\bf\tilde{r}}$.
Using the expressions for the positions and velocities in regions I and III and the components of ${\bf A^{(t)}}$, one may easily verify (41) from both sides. Working explicitly with vector components, it is easy to see how the net difference in the vector potential plays a crucial role in ensuring the equality here. Eqn. (41), which visibly demonstrates the need and presence of the difference in the vector potential, is another generalization of the velocity kick formula in (1), when we have a uniform magnetic field perpendicular to the plane of motion. When $\omega_{B}=0$, it reduces to the case without a magnetic field. In addition, as shown earlier, when both the magnetic field and the electric pulse are present, there is indeed a change in relative separation.
IV Conclusions